What is the molecular formula of the alkane? Repl

What is the molecular formula of the alkane?

What is the molecular formula of the alkane?  Replace the subscript letters with numbers. Formatting counts.
Answer #1
Concepts and reason

The concept used to solve this problem is C-13 rule.

To determine the molecular formula of a compound, C-13 rule is used. Firstly, a base formula is generated by dividing the molar mass of substance by 13. The numerator (n)\left( n \right) or quotient obtained is equal to the number of carbon atoms and the sum of numerator and remainder (r)\left( r \right) is equal to the number of hydrogen atoms in the base formula. The base formula comes out to be CnHn+r{{\rm{C}}_n}{{\rm{H}}_{n + r}} .

Fundamentals

A mass spectrum is a plot of relative abundance of ion vs m/Zm/Z (mass to charge ratio).

According to C-13 rule, the molecular formula of a compound is a multiple, nn of 13 with a remainder rr . The rule is named C-13 because the molar mass of CH{\rm{CH}} is 13. So, the molecular formula used for interpreting the mass spectra is CnHn+r{{\rm{C}}_n}{{\rm{H}}_{n + r}} . The formula is adjusted in case of heteroatoms. For instance, if the heteroatom O{\rm{O}} is present in the compound, O{\rm{O}} is added and CH4{\rm{C}}{{\rm{H}}_{\rm{4}}} is subtracted.

The molecular ion mass is 100.

The molecular formula of the compound can be calculated using C-13 rule as follows:

10013=7r9\frac{{100}}{{13}} = 7r9

That is,

n=7r=9\begin{array}{l}\\n = 7\\\\r = 9\\\end{array}

The base formula is CnHn+r{{\rm{C}}_n}{{\rm{H}}_{n + r}} .

Substitute 7 for nn and 9 for rr in the base formula to obtain the molecular formula.

Molecularformula=C7H7+9=C7H16\begin{array}{c}\\{\rm{Molecular formula}} = {{\rm{C}}_{\rm{7}}}{{\rm{H}}_{{\rm{7 + 9}}}}\\\\ = {{\rm{C}}_{\rm{7}}}{{\rm{H}}_{{\rm{16}}}}\\\end{array}

Therefore, the molecular formula of the alkane is C7H16{{\rm{C}}_{\rm{7}}}{{\rm{H}}_{{\rm{16}}}} . The compound is n-heptane.

Chemical Formula: C7H16

The mass fragmentation of the compound is as follows:

| Intz - ICO
mz=43
mi = 57
mz = 19
| nz = 71

This explains the signal at m/Z=29m/Z = 29 , m/Z=43m/Z = 43 , m/Z=57m/Z = 57 and m/Z=71m/Z = 71 .

The peak at m/Z=85m/Z = 85 is due to the removal of CH2{\rm{C}}{{\rm{H}}_{\rm{2}}} group from the parent ion.

Ans:

The molecular formula of the alkane is C7H16{{\rm{C}}_{\rm{7}}}{{\rm{H}}_{{\rm{16}}}} .